# SpectrumWorld: Artificial Intelligence Foundation for Spectroscopy

### v1

@@ 4 SpectrumLab
Model Signal Perception Semantic Generation Avg. TC QE FE ID GR EP PA PP SE FM MR FP IP DnG Perf. Closed-source MLLMs Claude-3.5-Sonnet 96.36 28.33 76.47 71.43 60.00 77.78 76.32 85.29 82.50 69.23 94.59 20.00 0 0 57.78 Claude-3.7-Sonnet 96.36 38.33 86.27 82.14 71.43 88.89 71.05 88.24 82.28 74.36 89.19 20.00 0 5.26 62.02 Claude-4-Sonnet 96.36 35.00 88.24 92.86 62.22 63.89 60.53 76.47 16.25 43.59 64.86 3.33 0 21.05 45.49 Claude-3.5-Haiku 94.55 31.67 50.98 92.86 66.67 75.00 76.32 76.47 67.50 64.10 81.08 10.00 0 0 52.71 Claude-4-Opus 96.36 33.33 86.27 92.86 73.33 83.33 71.05 85.29 32.50 76.92 86.49 16.67 0 5.26 54.15 GPT-4o 96.36 33.33 68.63 92.86 57.78 77.78 63.16 79.41 78.75 58.97 89.19 10.00 0 0 54.58 GPT-4.1 94.55 28.33 86.27 85.71 53.33 77.78 63.16 79.41 82.50 66.67 91.89 33.33 10.53 0 58.33 GPT-4-Vision 94.55 33.33 72.55 92.86 73.33 72.22 71.05 82.35 73.75 53.85 97.30 23.33 5.00 0 57.51 Grok-2-Vision 94.55 31.67 74.51 89.29 64.44 80.56 73.68 82.35 37.50 66.67 81.08 23.33 0 0 52.21 Qwen-VL-Max 94.55 36.67 90.20 92.86 60.00 80.56 78.95 88.24 32.50 71.79 91.89 43.33 0 5.26 56.64 Open-source MLLMs Qwen2.5-VL-32B-Instruct 92.73 26.67 37.25 71.43 57.78 44.44 31.58 61.76 0.00 5.13 45.95 20.00 0 0 30.75 Qwen2.5-VL-72B-Instruct 94.55 38.33 86.27 92.86 42.22 80.56 78.95 88.24 66.25 76.92 91.89 30.00 0 10.53 58.95 InternVL3-78B 96.36 38.33 70.59 71.43 48.49 75.00 81.58 88.24 62.50 69.23 83.78 23.33 0 5.26 55.19 Llama-3.2-11B-Vision-Instruct 34.55 11.67 13.73 25.00 20.00 41.67 15.79 29.41 7.50 5.13 21.62 0 0 0 14.26 Llama-3.2-90B-Vision-Instruct 38.18 10.00 35.29 25.00 17.78 27.78 28.95 20.59 21.25 5.13 43.24 0 0 0 18.78 DeepSeek-VL2 52.73 23.33 29.41 28.57 8.89 27.78 28.95 50.00 15.00 15.38 32.43 10.00 5.00 5.26 22.26 Doubao-1.5-Vision-Pro 98.18 33.33 78.43 92.86 66.67 83.33 68.42 88.24 67.50 56.41 89.19 6.67 0 0 55.65 Doubao-1.5-Vision-Pro-Thinking 96.36 35.00 78.43 67.86 53.33 80.56 73.68 91.18 68.75 66.67 91.89 66.67 5.00 5.26 61.01 Overall Avg. 86.35 29.81 67.21 75.60 53.21 68.83 61.29 74.51 49.71 52.56 75.98 23.63 1.42 3.51 49.54
@@ 5 Experiment
(1) Closed-source Models Lead Overall Performance with Claude-3.7-Sonnet Achieving Best Results. Claude-3.7-Sonnet emerges as the top-performing model with an overall average accuracy of 62.02%, securing top-2 scores in 7 out of 14 tasks. The model demonstrates exceptional capabilities across multiple dimensions: it leads in Quality Assessment (QE) with 38.33%, Elemental Compositional Prediction (EP) with 88.89%, and ranks second in several other tasks, including Spectrum Type Classification (TC) and Basic Property Prediction (PP). Closed-source models generally maintain a performance advantage, with the top 5 models by overall average all being proprietary solutions. However, this gap is narrowing, models like Doubao-1.5-Vision-Pro-Thinking (61.01%) and InternVL3-78B (55.19%) are approaching or even surpassing some closed-source counterparts in specific benchmarks.

### v3

@@ 4 SpectrumLab
Model Signal Perception Semantic Generation Avg. TC QE FE ID GR EP PA PP SE FM MR FP IP DnG Perf. Closed-source MLLMs Claude-3.5-Sonnet 96.36 28.33 76.47 71.43 60.00 77.78 76.32 85.29 82.50 69.23 94.59 20.00 0 0 57.78 Claude-3.7-Sonnet 96.36 38.33 86.27 82.14 71.43 88.89 71.05 88.24 82.28 74.36 89.19 20.00 0 5.26 62.02 Claude-4-Sonnet 96.36 35.00 88.24 92.86 62.22 63.89 60.53 76.47 16.25 43.59 64.86 3.33 0 21.05 45.49 Claude-3.5-Haiku 94.55 31.67 50.98 92.86 66.67 75.00 76.32 76.47 67.50 64.10 81.08 10.00 0 0 52.71 Claude-4-Opus 96.36 33.33 86.27 92.86 73.33 83.33 71.05 85.29 32.50 76.92 86.49 16.67 0 5.26 54.15 GPT-4o 96.36 33.33 68.63 92.86 57.78 77.78 63.16 79.41 78.75 58.97 89.19 10.00 0 0 54.58 GPT-4.1 94.55 28.33 86.27 85.71 53.33 77.78 63.16 79.41 82.50 66.67 91.89 33.33 10.53 0 58.33 GPT-4-Vision 94.55 33.33 72.55 92.86 73.33 72.22 71.05 82.35 73.75 53.85 97.30 23.33 5.00 0 57.51 Grok-2-Vision 94.55 31.67 74.51 89.29 64.44 80.56 73.68 82.35 37.50 66.67 81.08 23.33 0 0 52.21 Qwen-VL-Max 94.55 36.67 90.20 92.86 60.00 80.56 78.95 88.24 32.50 71.79 91.89 43.33 0 5.26 56.64 Doubao-1.5-Vision-Pro 98.18 33.33 78.43 92.86 66.67 83.33 68.42 88.24 67.50 56.41 89.19 6.67 0 0 55.65 Doubao-1.5-Vision-Pro-Thinking 96.36 35.00 78.43 67.86 53.33 80.56 73.68 91.18 68.75 66.67 91.89 66.67 5.00 5.26 61.01 Open-source MLLMs Qwen2.5-VL-32B-Instruct 92.73 26.67 37.25 71.43 57.78 44.44 31.58 61.76 0.00 5.13 45.95 20.00 0 0 30.75 Qwen2.5-VL-72B-Instruct 94.55 38.33 86.27 92.86 42.22 80.56 78.95 88.24 66.25 76.92 91.89 30.00 0 10.53 58.95 InternVL3-78B 96.36 38.33 70.59 71.43 48.49 75.00 81.58 88.24 62.50 69.23 83.78 23.33 0 5.26 55.19 Llama-3.2-11B-Vision-Instruct 34.55 11.67 13.73 25.00 20.00 41.67 15.79 29.41 7.50 5.13 21.62 0 0 0 14.26 Llama-3.2-90B-Vision-Instruct 38.18 10.00 35.29 25.00 17.78 27.78 28.95 20.59 21.25 5.13 43.24 0 0 0 18.78 DeepSeek-VL2 52.73 23.33 29.41 28.57 8.89 27.78 28.95 50.00 15.00 15.38 32.43 10.00 5.00 5.26 22.26 Overall Avg. 86.35 29.81 67.21 75.60 53.21 68.83 61.29 74.51 49.71 52.56 75.98 23.63 1.42 3.51 49.54
@@ 5 Experiment
(1) Closed-source Models Lead Overall Performance with Claude-3.7-Sonnet Achieving Best Results. Claude-3.7-Sonnet emerges as the top-performing model with an overall average accuracy of 62.02%, securing top-2 scores in 7 out of 14 tasks. The model demonstrates exceptional capabilities across multiple dimensions: it leads in Quality Assessment (QE) with 38.33%, Elemental Compositional Prediction (EP) with 88.89%, and ranks second in several other tasks, including Spectrum Type Classification (TC) and Basic Property Prediction (PP). Closed-source models generally maintain a performance advantage, with the top 5 models by overall average all being proprietary solutions. However, this gap is narrowing, models like Qwen2.5-VL-72B-Instruct(58.95%) and InternVL3-78B (55.19%) are approaching or even surpassing some closed-source counterparts in specific benchmarks.